### v1

# GRASP: A Grid-Based Benchmark for Evaluating Commonsense Spatial Reasoning

Abstract
Spatial reasoning, an important faculty of human cognition with many practical applications, is one of the core commonsense skills that is not purely language-based and, for satisfying (as opposed to optimal) solutions, requires some minimum degree of planning. Existing benchmarks of Commonsense Spatial Reasoning (CSR) tend to evaluate how Large Language Models (LLMs) interpret text-based spatial descriptions rather than directly evaluate a plan produced by the LLM in response to a spatial reasoning scenario. In this paper, we construct a large-scale benchmark called GRASP, which consists of 16,000 grid-based environments where the agent is tasked with an energy collection problem. These environments include 100 grid instances instantiated using each of the 160 different grid settings, involving five different energy distributions, two modes of agent starting position, and two distinct obstacle configurations, as well as three kinds of agent constraints. Using GRASP, we compare classic baseline approaches, such as random walk and greedy search methods, with advanced LLMs like GPT-3.5-Turbo and GPT-4o. The experimental results indicate that even these advanced LLMs struggle to consistently achieve satisfactory solutions.
1 Introduction
Spatial reasoning, a fundamental aspect of human cognition, is a core commonsense skill that allows people to navigate, interact with, and manipulate their environments effectively [8, 17, 11, 10]. This cognitive ability involves recognizing and understanding the spatial relations among different entities, [19], as well as actively using it to navigate through space while manipulating objects within it [15], which is essential in day-to-day activities such as walking and driving. An adequate spatial reasoning ability means more than just recognizing the relative spatial relationships between objects but also using such relationships, in combination with planning skills, to achieve satisfying, if not optimal, solutions for tasks at hand.
With the advancement of Large Language Models (LLMs) in understanding and generating natural language [1, 29, 24, 23], exhibiting near human-level commonsense reasoning abilities [12, 30, 20], as well as their increasing integration into applications requiring spatial awareness [9, 3, 27, 31], evaluating their spatial reasoning capabilities has raised significant attention [21, 25, 26, 2]. Several benchmarks, such as SpartQA [16], StepGame[22], and BabyAI[4], have been proposed to evaluate these abilities by simulating visual or textual spatial relationships. However, these benchmarks focus on interpreting spatial descriptions, whether textually or visually, and neglect the practical aspect of spatial reasoning.
The ability to use interpreted spatial information to navigate and interact with the environment in a commonsensical manner is termed Commonsense Spatial Reasoning (CSR) ability. It is essential to effectively navigate around environments. For example, an agent must understand the fundamental spatial principles such as ‘if an obstacle is in the way, an alternative path that goes around the obstacle should be considered’ and ‘if two resources are close together, it is more efficient to collect them in one trip rather than multiple trips.’ Such CSR enables the agent to make effective decisions with regard to the spatial layout, such as prioritizing the collection of nearby resources before exploring further and adapting to obstacles by rerouting effectively.
[…]
GRASP is a set of environments that consists of 2000 different grid instances and 8 distinct combinations of agent constraints, resulting in 16000 environments. Each grid is constructed as a two-dimensional array of cells, each of which can be empty, an obstacle, or contain a unit of energy. The objective for the agent is to collect as much energy as possible within a fixed number of steps, simulating real-world scenarios where agents must navigate through the environments and plan a path to collect resources. The environment includes various patterns of energy arrangement, such as random, vertically-skewed, horizontally-skewed, clustered, and spiral distributions. These patterns challenge the agent’s ability to recognize the differences and adapt its strategy based on different distributions of resources. Additionally, obstacles are randomly placed within the grid and the agent’s starting position is randomized within specified inner or outer regions of the grid, testing the agent’s adaptability to different starting scenarios. Furthermore, we introduce various constraints that affect the behaviors of the agent, starting with a limited set of actions that include basic movements and, in a more complex setup, additional diagonal movements. Additional Constraints, such as the maximum number of energy units the agent can carry and the cost associated with each step, are designed to further test the agent’s CSR and associated planning abilities.
2 Related Work
[…]
Additionally, BabyAI [4] provides a grid-based platform to support the research of including humans in the loop for grounded language learning. Minigrid [5] provides modular and customizable grid-based environments designed for goal-oriented tasks in the reinforcement learning paradigm. However, these benchmarks primarily focus on evaluating an agent’s ability to follow natural language instructions with spatial information, while our benchmark investigates the agent’s ability to autonomously navigate the grid and collect resources using spatial information. More importantly, they see the task as needing further model optimization, while we argue that spatial reasoning should be treated as a commonsense ability and evaluated accordingly [20]. Moreover, in our grid environment, the agent begins with complete knowledge of the grid’s layout and all necessary task information. Unlike other benchmarks that require agents to explore and actively gather data, our framework provides all information upfront, eliminating the need for sequential interaction with the environment. Table 1 provides a comparison between GRASP and other related benchmarks.
Table 1: GRASP in Comparison to Other Benchmarks
Benchmark Spatial Representation Planning Required? Optimization Needed? Information Scope Spatial Commonsense Needed?
Minigrid [5] visual yes yes partially no
BabyAI [4] visual yes yes partially no
PPNL [2] textual yes yes fully no
SPARTQA [16] textual no no fully no
StepGame [22] textual no no fully no
GRASP (ours) hybrid yes no fully yes
3 GRASP: Benchmark Construction
Figure 1: The pipeline for constructing one instance of the grid environment in the GRASP benchmark.
GRASP evaluates the commonsense spatial reasoning ability of LLMs by testing their capability to navigate in a grid-base environment while assigned an energy collection task. The benchmark is a combination of grid environments and agent constraints. The grid environment, $G$, can be understood as a 2D square array of cells $C_{ij}$, indexed from 0 to $k$ and from left to right and top to bottom, with $i$ representing row index and $j$ representing column index: $C_{00},...,C_{0k},C_{10},...,C_{1k},...,C_{k0},...,C_{kk}$. A cell can either be empty ($C_{ij}=Null$), contain one unit of energy ($C_{ij}=E$), have an obstacle ($C_{ij}=O$), or be the starting point of the agent ($C_{ij}=A$).
As for the agent constraints, the agent is allowed pre-defined ($n$) steps. At each step, the agent can choose from a set of movement-related actions ($\mu$) that allows it to move around the grid environment and a set of resource-related actions ($\rho=\{TAKE,DROP\}$) that enables it to take the energy from or drop the collected energy to the cell that it is in. The agent’s objective is to collect as much energy as possible in $n$ steps. In section 3.1, we introduce three different controls that can affect the layout of the grid environment. In section 3.2, we discuss three types of constraints that the agent must follow in accomplishing the task in our grid environment. In 3.3, we explain how such a grid environment is represented to an LLM-based agent. Figure 1 shows the entire pipeline for constructing one instance of the grid environment.
3.1 Grid Environments
We establish the side length $k$ of the grid environment to be 11. As a result, there are $11\times 1$ 11 = 121 cells in the grid. Based on this setup, we construct a set of grid templates that can be further instantiated to become a grid instance. Each grid template consists of three controls: Energy assignment, Obstacle construction, and Starting position. Each control can take several different values.
Specifically, for the energy assignment, the random distribution assigns an equal probability of having one unit of energy in each cell, setting a baseline test bed for the agent’s ability. The vertically-skewed and horizontally-skewed distributions introduce asymmetry in the energy assignment, testing for any spatial bias in the agent’s decisions. Furthermore, cluster distribution introduces localized energy clusters, and spiral distribution generates an energy path from the center, both exploring the agent’s ability to identify and target specific regions or parts of the grid. Details about the five distinct types of energy assignment are described in the following:
  • Random distribution of energy means that each cell in the grid has the same probability of containing one unit of energy. This probability is determined in during instantiation and follows a uniform distribution between 0.3 and 0.7. For one grid template, $P(C_{ij}=E)=p,p\sim U(0.3,0.7),\forall i,j$.
  • Vertically-skewed distribution means the cells in the top-half-part of the grid (rows 0 to 5) are more likely or less likely (determined in instantiation) to contain one unit of energy than the bottom-half-part of the grid (rows 6 to 10). In instantiation, the probability $p_{top}$ of a top-half-part cell containing one unit of energy is randomly instantiated from one of the two uniform distributions: $U(0.3,0.4)$ and $U(0.6,0.7)$. Then, the probability $p_{bottom}$ of a bottom-half-part cell containing one unit of energy is defined as $p_{bottom}=1-p_{top}$. Formally, $P(C_{ij}=E)=p_{top}$, $\forall$ $0<=i<=5,j$ and $P(C_{ij}=E)=p_{bottom}$, $\forall$ $6<=i<=10,j$. As for $p_{top}$,
    $p_{\text{top}}\sim\begin{cases}U(0.3,0.4)&\text{with probability }0.5\\ U(0.6,0.7)&\text{with probability }0.5\end{cases}$
  • Horizontally-skewed distribution is similar to the vertically-skewed distribution. According to this distribution, cells in the left-half-part of the grid (columns 0 to 5) are more likely or less likely (determined in instantiation) to contain one unit of energy than the right-half-part of the grid (columns 6 to 10). In instantiation, the probability $p_{left}$ of a left-half-part cell containing one unit of energy is randomly instantiated from one of the two uniform distributions: $U(0.3,0.4)$ and $U(0.6,0.7)$. Then, the probability $p_{right}$ of a right-half-part cell containing one unit of energy is defined as $p_{right}=1-p_{left}$. Formally, $P(C_{ij}=E)=p_{left}$, $\forall$ $i,0<=j<=5$ and $P(C_{ij}=E)=p_{right}$, $\forall$ $i,6<=j<=10$. As for $p_{left}$,
    $p_{\text{left}}\sim\begin{cases}U(0.3,0.4)&\text{with probability }0.5\\ U(0.6,0.7)&\text{with probability }0.5\end{cases}$
  • Cluster distribution creates a number of energy clusters around the grid. Each cluster enables the cells in a 3 by 3 grid to be filled with energy. The centers of the clusters are chosen randomly. Hence, it is possible for clusters to overlap each other. The number of clusters is randomly chosen between 3, 4, and 5 during instantiation. All other cells do not contain any energy. Formally, if the center of the cluster is denoted by the cell $C_{ab}$, then $P(C_{ij}=E)=1$, $\forall$ $a-1<=i<=a+1,b-1<=j<=b+1$ and $a\sim U\{0,10\}$ and $b\sim U\{0,10\}$.
  • Spiral distribution creates a spiral-like energy trail starting from the center cell (the cell in row 5, column 5). The spiral progresses outward with variations in the angle and radius. The angle is adjusted with random variation, and the radius also varies with random fluctuations. The angle $\theta$ at step $i$ is given by: $\theta=\frac{i}{10}+\epsilon_{\theta}$. $\epsilon_{\theta}$ is a random variation in the angle and $\epsilon_{\theta}\ \sim$ $U(-0.2,0.2)$. The radius $r$ at step $i$ is given by: $r=\frac{i}{11\cdot 10/(2\pi)}+\epsilon_{r}$, where $\epsilon_{r}$ is a random variation in the radius and $\epsilon_{r}\ \sim$ $U(-0.2,0.2)$. The coordinates $(x,y)$ of each cell are then calculated as: $x=\text{int}\left(5+r\cdot\cos(\theta)\right)$, $y=\text{int}\left(5+r\cdot\sin(\theta)\right)$. The probability of a cell $(x,y)$ containing one unit of energy is 1 if it lies on the spiral path defined by the above functions and 0 if not.
In the instantiation process, each grid is assigned one specific energy distribution type. Each cell in this grid is constructed to have one unit of energy according to its associated probability that is described above.
In addition to energy, a grid template can also contain obstacles. If a grid is established to have obstacles, each cell of the grid has a 0.1 probability of being an obstacle. That is, $P(C_{ij}=O)=0.1,\forall i,j$. If a cell contains an obstacle, the agent is not allowed to cross the cell and take any energy from the cell. Finally, the agent’s starting position is determined. There are two positions: inner position and outer position. The inner position is defined as any cell in the square formed by rows 3 and 7 and columns 3 and 7 (i.e., $C_{ij},\forall$ $73<=i<=7,3<=j<=$ 7). The outer position is defined as any cell in the grid that is not in the inner position (i.e., $C_{ij},\forall$ $20<=i<=$ 2 or $108<=i<=$ 10, and $20<=j<=$ 2 or $108<=j<=$ 10). In instantiation, the agent’s starting position configuration is randomly chosen between the inner position and outer position. Then, a specific cell in the chosen position configuration is selected uniformly at random as the starting position, ignoring any energy or obstacle that is already in that cell.
In total, we generated 100 instances for each combination of the five different energy distributions, two distinct obstacle configurations, and two modes of agent starting positions, resulting in $100\times 5\times 2\times 2=2000$ grid instances.
3.2 Agent Constraints
Along with different grid environments, we also consider constraints that the agent needs to follow when accomplishing the task. First of all, we fix the number of steps the agent can take to be $n=20$ steps. Additionally, the agent cannot move beyond the boundary of the grid, and any invalid action will result in no change to the environment.
For each step, we design two sets of movement-related actions that the agent is allowed to take: $\mu_{1}=$UP, DOWN, LEFT, RIGHT}, $\mu_{2}=$UP, DOWN, LEFT, RIGHT, UPLEFT, UPRIGHT, DOWNLEFT, DOWNRIGHT}. The outcome of each action corresponds to the characteristics suggested by its name. For example, UP allows you to move one cell up in one step, and UPLEFT allows you to move diagonally one cell up and left in one step. The two sets of movement-related actions offer the agent varying degrees of movement flexibility, with one offering basic movements while the other allows for additional diagonal movements. Additionally, we design two resource-related actions: $\rho=\{TAKE,DROP\}$. TAKE allows the agent to take the energy from the cell if the agent is moved to a cell with energy, and DROP allows the agent to drop all the energy that the agent carries to the cell that the agent is currently on.
Furthermore, we design two constraints that could potentially affect the agent’s behaviors. The first constraint, if active, adds a limit on the number of energy units that the agent can carry to a maximum of 2 units, requiring the agent to drop energy before acquiring any more of it beyond two units. The second constraint, if active, adds a cost to each action, forcing the agent to consider the energy expenditure of its actions. This cost is defined as 0.3 units of energy per action. Combining each of the 2000 grid environment instances with the three different agent constraints, we have a total of $2000\times 2\times 2\times 2=16000$ different instances of the benchmark.
3.3 Grid Representation to An LLM-based Agent
So far, we have described the grid as an abstract environment. However, to present it to an LLM-based agent, the grid needs a concrete representation. Previous works either use text descriptions [16, 22] or involve another modality (i.e., visual) [4, 5]. While textual description can convey spatial information without visual elements, it comes with several disadvantages. For example, spatial relationships can be described in multiple ways, leading to bias in the way of description. Moreover, there is a limitation on the granularity of description, as it is unclear at what level should the spatial relationships be detailed. Describing complex environments can also result in excessive information, making it challenging to provide an accurate depiction. In contrast, visual representations offer detailed spatial information about the environment. However, they also present their own challenges, as the completion of tasks depends on the agent’s visual processing abilities, which can have inherent flaws. Moreover, visual representations can be overly complex, potentially confusing the agent
Our approach balances the benefits and drawbacks of both text description and visual representations by considering a hybrid representation. We present the grid as an 11 by 11 matrix labeled with coordinates from 0 to 10 on both the x-axis (top) and y-axis (left). The grid is constructed with cells separated by dashed lines, forming a series of rows and columns. The cells are separated vertically using “+ - - - +" and horizontally by “|". Each cell contains a character that represents if there is a unit of energy, an obstacle, or an agent on it. Specifically, a unit of energy is represented as “E”; an obstacle is represented as “O”; the agent is represented as “A”. If the cell contains none of the three entities, it will be left as a whitespace character with length one. Additionally, to align the cells both horizontally and vertically, an extra whitespace character is added on both sides of the cell content. An example of the grid is shown in Figure 2, and its text form is shown in Appendix A.1.
Our approach eliminates the need for textual descriptions by actively representing the spatial relationships within text space. It has been shown that LLMs are capable of understanding such representation [18]. Furthermore, there is no additional visual processing capability requirement for LLMs, allowing the evaluation to concentrate on their ability to reason using explicit spatial information rather than their ability to interpret visual information. Our approach enhances the experimental control over biases and hallucinations.
4 Experiments
Two baseline approaches are evaluated. The first is the random walk method. This method simulates an agent who has no knowledge about the grid, the agent constraints regarding the limit on the number of energy that the agent can carry, and the cost of each action, but is only aware of the set of available actions. The random walk method randomly selects one out of the four or eight movement-related actions, depending on the available actions. Following each of the movement-related actions, this approach will select the TAKE action to take the energy from the cell, whether or not the cell has any energy. This ‘move and take’ behavior will continue for 6 times. Then, this approach will select the complement of the movement-related actions that were selected before in reverse order to get back to the starting position and finally select DROP to drop all the energy. The complement of a movement-related action is defined as the complementary action that exactly reverses the movement. For example, if the initial movement-related action was ‘UP,’ the complementary action would be ‘DOWN.’ In total, this will take 19 steps, which is within the required 20 steps.
The second baseline approach is the greedy search method. This method simulates an agent with perfect knowledge of the grid environment but no planning abilities and no knowledge about the agent constraints except the available movement-related actions. This agent starts from the agent’s current position and uses a Breadth-First Search algorithm to navigate the grid. It considers all four $\mu_{1}$ or eight $\mu_{2}$ possible movement-related actions, depending on the available actions. At each step, it checks the neighboring cells in all directions. The sequence of the neighboring cells in consideration is randomized. If a neighboring cell contains energy, the search stops, and the path to this cell is recorded. If all neighboring cells are either empty or obstacles, the search continues to the next set of neighboring cells. Once the nearest energy cell is found, the recorded path is used to move the agent step-by-step towards that cell. This will continue until a neighboring cell that has energy is too far from the current position, and the remaining number of steps is not enough for the agent to go there and get back to the starting position using the complement of past movement-related actions in reverse order. At this point, the agent simply uses the complement of past movement-related actions in reverse order to get back to the starting cell and drop all the collected energy. The pseudo-code of both methods is provided in Appendix A.2.
Additionally, we evaluate the performance of two LLMs: GPT-3.5-Turbo (i.e., gpt-3.5-turbo-0125) and GPT-4o. All models are used with a temperature setting of 0 to ensure reproducibility. We only test these models with the basic zero-shot prompting method. To construct the prompt used for the LLMs, we utilize all available information, both from the grid environment and agent constraints. Instructions and agent constraints are provided to LLMs in the system prompt, and the grid is provided in the user prompt; the LLMs are required to provide a full list of actions in the response, with no further interaction with the environment. The specific prompts used are detailed in Appendix A.3. The experiments are conducted using one-tenth of the entire GRASP benchmark, for grid instances indexed from 0 to 9. Associated results are reported below.
4.1 Results
Table 2: The average number of steps taken by the agents (displayed in the Length column) and the average units of energy collected and dropped to the starting position by the agents (displayed in the Energy column), are calculated for each control setting. The green numbers in each row represent the maximum energy collected by every agent, while the red numbers represent the minimum ones.
$\downarrow$ Control &\&& its Values $\searrow$ Agent $\rightarrow$ Random Walk Greedy Search GPT-3.5- Turbo GPT-4o
Length Energy Length Energy Length Energy Length Energy
Energy Distribution Random 19.00 -1.26 18.75 -0.14 19.70 -1.96 16.60 -1.20
Vertically-skewed 19.00 -1.43 18.65 -0.23 19.59 -1.66 17.40 -1.55
Horizontally-skewed 19.00 -1.33 18.62 -0.07 19.74 -2.12 17.28 -1.43
Cluster 19.00 -1.90 18.81 -0.09 18.58 -2.09 16.65 -1.28
Spiral 19.00 -1.77 18.73 -0.15 19.41 -2.34 17.09 -1.45
Obstacle Yes 19.00 -1.57 18.70 -0.19 19.33 -2.09 17.18 -1.58
No 19.00 -1.51 18.72 -0.09 19.48 -1.98 16.83 -1.18
Starting Position Inner Position 19.00 -1.50 18.73 -0.02 19.38 -1.92 17.04 -1.34
Outer Position 19.00 -1.58 18.70 -0.25 19.43 -2.14 16.96 -1.42
Movement-related Action Set $\mu_{1}$ 19.00 -1.21 18.54 0.80 19.12 -1.89 16.75 -1.13
$\mu_{2}$ 19.00 -1.87 18.89 -1.07 19.69 -2.17 17.26 -1.63
Energy Carrying Limit No Limit 19.00 -1.40 18.73 0.83 19.78 -2.00 16.72 -1.08
2 Units 19.00 -1.68 18.69 -1.10 19.03 -2.06 17.28 -1.68
Energy Cost Per Step 0 Unit 19.00 1.30 18.72 2.66 19.40 1.24 17.15 1.16
0.3 Unit 19.00 -4.38 18.70 -2.93 19.41 -5.30 16.86 -3.92
Average 19.00 -1.54 18.71 -0.14 19.41 -2.03 17.00 -1.38
Table 2 displays the average number of steps taken and units of energy collected and dropped for the two baseline agents (Random Walk and Greedy Search) and the two LLM-based agents (GPT-3.5-Turbo and GPT-4o), evaluated using GRASP. Most Energy results are negative numbers because they are averaged across all control settings, including a 0.3 unit energy cost per step that can lead to a negative energy result if the energy cost exceeds the energy collected. When considering the overall performance of each agent, we find that the Greedy Search agent performs the best across all settings, while GPT-3.5-Turbo performs the worst in 13 out of 14 settings. Additionally, GPT-4o performs better than GPT-3.5-Turbo but worse than the Greedy Search agent and is roughly on par with the Random Walk agent.
Focusing on the performance within each control, we observed Energy differences of 0.4, 0.5, and 0.6 between the two settings of ‘Obstacle,’ ‘Movement-related Action Set,’ and ‘Energy Carrying Limit’ for GPT-4o. This agent performs the best without any obstacles and energy carrying limit while using the action set $\mu_{1}$, which only includes the four regular movements (i.e., UP, DOWN, LEFT, RIGHT) without the diagonal movements. However, these differences are less apparent for GPT-3.5-Turbo, with the maximum difference being only 0.56 units of energy in the ‘Energy Carrying Limit’ control. Notably, GPT-4o achieves, on average, 5.08 units less energy when given the constraint regarding the energy cost per step than when given no such constraint. A similar phenomenon is also observed for the GPT-3.5-Turbo agent. Additionally, the ‘Energy Distribution’ and ‘Starting Position’ controls seem to play a lesser role in affecting the performance of GPT-4o and GPT-3.5-Turbo, as the difference in the amount of energy collected and dropped is less apparent (with the maximum difference being less than 0.35). However, for GPT-3.5-Turbo, the ‘Energy Distribution’ control shows a more significant impact, with a maximum difference of 0.68 between the ‘Spiral’ and ‘Vertically-skewed’ distributions. Turning to the average length of the steps taken by each agent, we observe that GPT-4o takes the least number of steps among the four agents, often between 16 to 17 steps, while GPT-3.5-Turbo takes the most number of steps, taking 19 to 20 steps.
Figure 2: The path formed by the actions of the four agents (Random Walk, Greedy Search, GPT-3.5-Turbo, and GPT-4o) in the shown grid environment. The agents are under the following constraints: unlimited energy carrying limit, zero energy cost per step, and the $\mu_{1}$ movement-related action set (i.e., UP, DOWN, LEFT, RIGHT). Arrows indicate the movement-related actions, and a yellow star next to each cell signifies the agents’ ‘TAKE’ action in that cell.
Figure 2 illustrates a typical action path of the four agents within the shown instance of the grid environment. We observe that the Random Walk agent wanders through the grid without a clear objective and gives the ‘TAKE’ action in cells without energy, while the Greedy Search agent exhibits a more object-oriented behavior, moving directly towards cells with energy and taking it. Focusing on the LLM-based agent, we find that GPT-3.5-Turbo includes cells with energy in its path but often fails to collect energy, repeatedly attempts to move through the cell with an obstacle, and does not return to the starting cell to drop the collected energy. In contrast, the GPT-4o agent demonstrates more intelligent behavior by actively going towards cells with energy and collecting them while avoiding obstacles, yet it still shows inefficient behaviors, such as missing nearby energy cells with energy on them and taking extra and unnecessary steps to achieve the same objectives. These behaviors also reflect the performance of each agent observed in Table 2.
5 Discussion
In this paper, we introduce GRASP, a comprehensive benchmark designed to evaluate the Commonsense Spatial Reasoning (CSR) abilities of LLMs within structured grid environments. GRASP assesses agents’ abilities to navigate the grid environments and collect energy, presenting 16,000 diverse environments characterized by five types of energy distributions, two distinct obstacle configurations, two modes of agent starting position, and three kinds of agent constraints, including two sets of movement-related actions, two limits on the amount of energy that can be carried, and two different energy costs for each step. We evaluate the performance of two classic algorithms (Random Walk and Greedy Search) and advanced LLMs (GPT-3.5-Turbo and GPT-4o) using GRASP. Our findings show that GPT-3.5-Turbo under-performs compared to the Random Walk agent, often unable to successfully navigate the grid, avoid obstacles, and return to the starting point to drop the collected energy, suggesting a lack of CSR abilities. On the other hand, GPT-4o performs similarly to the Random Walk agent but is still less efficient than the Greedy Search agent. While GPT-4o can recognize spatial relationships and navigate around the grid environments, showing a primitive capability of CSR, its use of spatial information to come up with efficient plans remains limited.
Our study comes with several limitations. Firstly, the synthetic nature of GRASP’s grid environments may not fully capture the complexity of real-world CSR tasks, which include more dynamic and unpredictable environments with a greater variety of objects and incomplete or even hidden spatial information. Furthermore, GRASP focuses on commonsense spatial reasoning but does not consider the multi-modal nature of many real-world tasks where abundant visual information is available. Despite the current limitations and potential hallucination problems in LLMs’ multi-modal abilities, a reliable visual interpretation capability could provide the models with a richer set of spatial information than text alone. Additionally, we only consider scenarios with a single agent, whereas real-world environments often involve multiple agents, requiring more sophisticated behaviors, such as interacting and planning together with the other agents. Furthermore, our evaluation was limited to zero-shot prompting methods for LLMs, potentially underestimating their true capabilities.
To address these limitations, GRASP should be expanded to include more dynamic environments that better simulate real-world CSR tasks, preferably developed using real-world data instead of pure synthetic ones. Additionally, with the rapid advancement of multi-modal foundational models, it is crucial to integrate datasets that combine visual and textual spatial information to better capture the complexity of the CSR tasks and provide a more comprehensive evaluation of the models. Extending the benchmark to scenarios involving multiple agents would also more accurately reflect the intricacies of real-world environments. Furthermore, investigating more sophisticated prompting methods could improve the performance of LLMs in these tasks, providing a more holistic evaluation of their commonsense spatial reasoning abilities.
In conclusion, we have presented a novel benchmark for understanding and evaluating the commonsense spatial reasoning abilities of LLMs, albeit with few limitations. Our findings highlight the current limitations of these models in efficiently completing commonsense spatial reasoning tasks and underscore the need for continued research and development. By providing a structured and diverse benchmark, GRASP lays the foundation for future advancements in LLM capabilities, ultimately contributing to the development of more robust and intelligent systems capable of performing a wide range of commonsense spatial reasoning tasks.